OpenStudy (anonymous):

All of the following expressions are equivalent except _____. 2 - x x - 2 -2 + x x + (-2)

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

Let's say 'x' is 3. 2 - 3 = ? 3 - 2 = ? -2 + 3 = ? 3 + (-2) = ?

OpenStudy (anonymous):

@Shikatanori Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Oppss sorry @iGreen.

OpenStudy (anonymous):

im a her not a his and i need the answer!

OpenStudy (anonymous):

yeah substitute anything you want :) on x

OpenStudy (anonymous):

Subtract/Add all those I gave you

OpenStudy (anonymous):

See which one is different fromt he rest.

OpenStudy (anonymous):

*the

OpenStudy (anonymous):

its D right?!

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

C?!

OpenStudy (anonymous):

Hahaahahha LoL can you Add anything?

OpenStudy (anonymous):

no im just guessing!

OpenStudy (anonymous):

dont ask me to do sum because 9 times outta 10 im gonna guess!

OpenStudy (anonymous):

Hahaha Ok ok .. let us change the x = 1 subtitute it its much more better :)

OpenStudy (anonymous):

a. 2-1=? b. 1-2=? c. -2+1=? d. 1+(-2)=?

OpenStudy (anonymous):

its freakin b

OpenStudy (anonymous):

?!?! ya'll stress me out!